Circular No. 6864

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

CM DRACONIS E. Guinan, Villanova University; D. Bradstreet, Eastern College; I.

Ribas, University of Barcelona; M. Wolf, Charles University; and G. Mc-Cook, Villanova University, report the detection of a probable brown dwarf or massive planet in the CM Dra (dM4.5+dM4.5) eclipsing-binary system: “Possible planetary transit eclipses (with depths of 0.08 mag) have been reported previously for CM Dra (see IAUC 6423, 6425). From extensive photometry obtained during Mar. 1996–Mar. 1998, seventeen eclipse timings were obtained. Least-squares solutions of the timing residuals show a sinusoidal variation in the arrival times of the eclipses, corresponding to an amplitude of 18 ± 2 s and an orbital period of ∼ 70.3 ± 1.5 days. Adopting the systemic mass of CM Dra (M_{1} + M_{2} = 0.448 M_{⊙}) and assuming a circular orbit, the size of the third body’s orbit is ∼ 0.27 AU. The mass of the third body is M_{3} ∼ 0.061 ± 0.004 M_{⊙} if its orbital plane is coplanar (or nearly so) with the eclipsing pair. The phasing of the light-time effect indicates that the object producing the effect is not the object that produced the planetary transit dimming events, so it is likely that CM Dra contains more than one substellar object. We are searching for possible transit eclipses from this body at the expected conjunction times; however, for a transit eclipse to occur, its orbital plane would have to be within 0^{o}.4 of exactly coplanar.”

GRB 980329 S. Klose, H. Meusinger, and H. Lehmann, Thüringer Landessternwarte

Tautenburg, report: “I-band images obtained on Mar. 29.8–30.0 UT, using the Tautenburg Schmidt telescope (+ Schmidt-focus CCD camera), show an object with magnitude I ∼ 20 at the position of the fading radio

source reported by Taylor et al. (http://gcn.gsfc.nasa.gov/gcn/gcn3/

040.gcn3; α = 7^{h}02^{m}38^{s}.0, δ = +38^{o}50′44′′, ± 1′′, equinox 2000.0). This object was no longer visible on I-band images taken on Mar. 31.8–31.9 with the Tautenburg telescope (I > 21); it was also not found on an Iband image taken on Apr. 1.85 with the Calar Alto 2.2-m telescope (+ CAFOS; I > 21). No R-band counterpart of this object is seen on Tautenburg images taken on Mar. 29.8–30.0 (R > 22). Images are available via

http://www.tls-tautenburg.de/research/grb.html.”

1998 April 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green